OpenStudy (anonymous):

if i know the lengths of two sides of a right triangle, how do i find the third?

OpenStudy (anonymous):

180 - tha sum of the two lengths

OpenStudy (anonymous):

so how do i find

OpenStudy (anonymous):

e third side ?

OpenStudy (anonymous):

*the third side

OpenStudy (anonymous):

pythagoras' theorem.

OpenStudy (anonymous):

can you do an example?

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Pythagoras%27_theorem

OpenStudy (anonymous):

ex. u have 24 degrees and 50 degrees, add them. so 74 is the answer.. subtract 74 from 180.

OpenStudy (anonymous):

thank you!!

OpenStudy (anonymous):

derek, what you're saying is completely wrong...

OpenStudy (anonymous):

.. crap. then how do you do it?

OpenStudy (anonymous):

well for starters you don't pick out random numbers like derek has done, given that you've said nothing about the triangle except that it has a right angle.

OpenStudy (anonymous):

how do i find the third side of a right triangle?

OpenStudy (anonymous):

nope. 180 is for the internal sum of a triangle. so we subtract the sum of the two sides

OpenStudy (anonymous):

go to this page: http://en.wikipedia.org/wiki/Pythagoras%27_theorem read the line which says that a^2 + b^2 = c^2. Look at the image of the triangle on the right where it marks which sides are a, b, and c. You know two of those side lengths. Now you can rearrange it to get the third. Use square root and whatever else.

OpenStudy (anonymous):

180 degrees is the sum of the angles of a triangle. bfowler wants to know the third length of a triangle, given two lengths.

OpenStudy (anonymous):

yup, i know. that's why we subtracted the sum of the two sides.

OpenStudy (anonymous):

derek, the things you are saying are complete gibberish. Hopefully you know that and are just trolling. You know you can't determine the two other angles of a triangle given only one angle, right? You can say what they sum to, but you can't say "oh yeah the sum is 74, so subtract 74 from 180" like you did up there. That's complete nonsense that you've pulled from nowhere.

OpenStudy (anonymous):

do you mind helping me out? i used the numbers 7 and 10. 7 is a and 10 is b. i squared them and got 49 and 100. no i have 49+100= c sq

OpenStudy (anonymous):

i am very sure of my answer,, anyways, it's up to her to decide. Peace!!! besides it's very late

OpenStudy (anonymous):

what do i do now

OpenStudy (anonymous):

so c^2 = a^2 + b^2 = 49 + 100 = 149. Take the square root of both sides. So c = sqrt(149)

OpenStudy (anonymous):

22201??

OpenStudy (anonymous):

i did 149 and then sq it

OpenStudy (anonymous):

Derek I see you were just stating an example where you could calculate the other angle if you knew two of them. That would be right if the question was to find the third angle, but it's to find the uknown third length.

OpenStudy (anonymous):

flutter,, the hell, that's not geometry. xD

OpenStudy (anonymous):

hahaha,, can't type f*ck?? turns to flutter, hahah??

OpenStudy (anonymous):

sqrt means to take the square root. You want to find a number x so that x * x = 149. It'll be a little larger than 12 since 12*12 = 144. Type sqrt(149) into google or something to see what it is.

OpenStudy (phi):

See http://www.khanacademy.org/math/algebra/ck12-algebra-1/v/pythagorean-theorem a short helpful video

OpenStudy (anonymous):

it says 22201.. now what

OpenStudy (anonymous):

You're squaring it, but you need the square root.

OpenStudy (anonymous):

The square root should be smaller than 149

OpenStudy (anonymous):

12.20655562

OpenStudy (anonymous):

correct.

OpenStudy (anonymous):

Correct, so that is the length of your third side:)

OpenStudy (anonymous):

now what

OpenStudy (anonymous):

so thats the answer

OpenStudy (anonymous):

Yep :)

OpenStudy (anonymous):

should you round it??

OpenStudy (anonymous):

sqrt(149) is an irrational number. It has an infinite number of decimal places. You can't avoid it.

OpenStudy (anonymous):

thank you guys so much.